X Company makes two products, A and B, and uses an activity-based costing overhead allocation system, with three cost pools and three cost drivers. Budgeted costs and driver information for 2020 were as follows:

Cost Pool 1 Cost Pool 2 Cost Pool 3
Budgeted costs $109,000     $54,000     $29,000    
Cost Drivers
  Product A 83,200     57,000     69,700    
  Product B 77,000     50,200     50,400    


What will the allocation to Product B be? [round overhead rate(s) to two decimal places].
Answer #1

Ans:

Overhead rate = cost / cost drivers

1 = 109000/(83200+77000) = $0.68

2 = 54000 /(57000+50200) = $0.50

3 = 29000 /(69700+50400) = $0.24

Allocation to product B = 77000*0.68 +50200*0.50 + 50400*0.24

= $89556